Ruby is thrice as old as Jade. Jade is 4 years older than Pearl. In ten years Ruby will be 2 years more than twice as old as Jade will then be. How old are they now?
Answer by Maths68(1474)

Ruby is thrice as old as Jade. Jade is 4 years older than Pearl. In ten years Ruby will be 2 years more than twice as old as Jade will then be. How old are they now?
Let
Present age of Ruby = r
Present age of Jade = j
Present age of Pearl = p


Ruby is thrice as old as Jade
r=3j............(1)
Jade is 4 years older than Pearl
j=p+4...........(2)

In ten years
Ruby will be = r+10
Jade will be = j+10

Then


Ruby will be 2 years more than twice as old as Jade will be then
r+10=2(j+10)+2
r+10=2j+20+2
r+10=2j+22
r=2j+22-10
r=2j+12.............(3)
Put the value of r from (1) to (3)
3j=2j+12
3j-2j=12
j=12
Put the value of j in (2)
j=p+4...........(2)
12=p+4
p+4=12
p=12-4
p=8
Put the value of j in (1)
r=3j
r=3(12)
r=36


Present age of Ruby = r = 36 years old
Present age of Jade = j = 12 years old
Present age of Pearl = p = 8 years old